OpenStudy (anonymous):

What is the value of the function v=3/2t - 5/2 when t = 5? ___?

OpenStudy (ccswims):

plug in 5 for 't' v = 3/2(5) - 5/2 v = 7.5 - 2.5

OpenStudy (anonymous):

5?

OpenStudy (ccswims):

yes :)